# Solving the SU($N$) Orbital Hatsugai-Komoto Model

### v2

@@ III.1 Mott insulator at $\langle\hat{n}\rangle=1$
$\mathcal{D}=\frac{1}{2nV}\sum_{\alpha,\mathbf{k}\in{\rm rBZ},\sigma\neq\sigma^{\prime}}\langle n_{\alpha\mathbf{k}\sigma}n_{\alpha\mathbf{k}\sigma^{\prime}}\rangle,$ (3)
@@
where $V$ is the number of sites in rBZ. This definition differs from that used in the SU($N$) Hubbard model. However, in both cases, $\mathcal{D}$ is proportional to the average interaction energy and the first derivative of the free energy with respect to $U$ thereby preserving this physical correspondence. In Fig. 1(e,f), the solid lines are the OHK results while the markers denote DQMC-Hubbard results. The double occupancy from 4-OHK and DQMC-Hubbard agrees quantitatively. Moreover, 4-OHK shows consistent results at higher density $\langle\hat{n}\rangle>1$, whereas DQMC data becomes oscillatory due to the severe fermionic sign problem. As previously reported for SU($N$) Hubbard Ibarra-García-Padilla et al. (2021), the sharp jump in $\partial\mathcal{D}/\partial\langle\hat{n}\rangle$ serves as a strong signature of the Mott insulating state persisting even in the high-temperature regime.
@@
From this benchmarking, we find that the 4-OHK model successfully captures the high-temperature $N$-dependent physics in the SU($N$) Hubbard model. Aside from the overestimate of the Mott gap, we find quantitative agreement between 4-OHK and DQMC-Hubbard results for the density $\langle\hat{n}\rangle$-dependence of physical quantities, such as $\mathcal{D}$. Additionally, the simplicity of 4-OHK makes it advantageous for exploring higher densities and lower temperatures (next section), where DQMC simulations are limited by the fermionic sign problem. These features establish OHK as a powerful tool for studying SU($N$) Hubbard physics beyond the reach of conventional numerical methods.
@@
Low temperature. In Fig. 2(a), we calculate the $\langle\hat{n}\rangle$ vs $\mu$ relation around $\langle\hat{n}\rangle=1$ at $U=8t$ and inverse temperature $\beta=200t^{-1}$ for the 4-OHK model with varying $N=2,3,4$. As expected from the high-temperature results in Fig. 1(a,b), we observe a sharp characteristic plateau, signalling the presence of a Mott charge gap. However, a surprising feature emerges: the Mott gap does not shrink monotonically with increasing $N$ despite appearing to do so at high temperatures in both 4-OHK (Fig. 1(a)) or DQMC simulations of the Hubbard model (Fig. 1(b,f)). Fig. 2(a) reveals the ordering $\Delta(N=3)>\Delta(N=4)>\Delta(N=2)$ for $U=8t$, deviating from the high-temperature behavior. This discrepancy arises because, at high temperatures, the number of excited states increases monotonically with $N$, leading to more pronounced thermal softening of the gap for larger $N$. By leveraging the 4-OHK model, we can explore both low- and high-temperature regimes, allowing us to disentangle intrinsic ground-state properties from thermal effects associated with excited states.
@@
Further, we calculate the charge gap at $\langle\hat{n}\rangle=1$ as a function of $U$ for various $N$ in the 4- and 8 (4$\times$2)-OHK models at zero temperature, as shown in Fig. 2(b) and (c), respectively. In the SU(2) case, a charge gap opens for any $U>0$ due to Fermi surface nesting at exactly $\langle\hat{n}_{\sigma}\rangle=1/2$ with ${\bf q}=(\pi,\pi)$ Feng et al. (2023); Ibarra-García-Padilla et al. (2023); Mai et al. (2024). For $N>2$, such a nesting is absent at $\langle\hat{n}\rangle=1$, where $\langle\hat{n}_{\sigma}\rangle=1/N<1/2$. Consequently, a finite critical $U_{c}$ is required to open the gap, as indicated in Fig. 2(b). Moreover, we find that $U_{c}(N=3)>U_{c}(N=4)$, likely due to the particular Fermi surface geometry at $U=0$. This trend, $U_{c}(N=3)>U_{c}(N=4)$ is also observed in the 8-OHK (Fig. 2) with only quantitative modification. Once the gaps are open for both SU($3$) and SU($4$) cases, we observe $\Delta(N=3)>\Delta(N=4)$, consistent with Fig. 2(a). Extending our calculation to larger $N$, we find in Fig. 2(b) the results collapse for $N>n$. This can be understood from the structure of the OHK model, which consists of independent $n$-site Hubbard clusters at each $\mathbf{k}$ in the rBZ. When a charge gap opens at $\langle\hat{n}\rangle=1$, these $n$-site clusters are filled with exactly $n$ particles each. When $N>n$, these particles occupy different SU($N$) flavors to minimize the kinectic energy. Further increasing $N$ does change this configuration, meaning that $n$-OHK would not be able to capture any additional $N$-dependent physics beyond $N=n+1$. For this reason, for the band HK and 2-OHK, the $\langle\hat{n}\rangle=1$ results overlap for all SU($N>1$) and SU($N>2$), respectively (see supplement). It is important to note that this phenomenon is specific to the filling $\langle\hat{n}\rangle=1$. As we will discuss in the next nection, near half-filling $\langle\hat{n}\rangle=N/2$ when the charge gap opens, each Hubbard cluster in OHK contains $Nn/2$ particles. In this case, $n$-OHK retains the non-trivial SU($N$) physics for any $N$ as long as $n\geq 2$.
@@
Figure 3: Charge gap at hall-filling ($\langle\hat{n}\rangle=N/2$) as a function of interaction strength $U$ for SU($N$) 4-OHK model.
III.2 Mott insulator at $\langle\hat{n}\rangle=N/2$
As mentioned above, another generalization of the half-filled SU($2$) Mott insulator is the SU($N$) Mott insulator at half-filling $\langle\hat{n}\rangle=N/2$ for even $N$ (no half-filled charge gap exists for odd $N$). At this filling, the non-interacting Fermi surface for each flavor remains the same as in the SU($2$) case, thereby preserving the effect of Fermi surface nesting. Hence, we expect the charge gap opens for any finite $U>0$. This expectation is confirmed in Fig. 3 where we present the half-filled charge gap as a function of $U$ for the SU(N) 4-OHK model. For a fixed interaction strength, the charge gap decreases monotonically with increasing $N$. As the number of flavors increases, more electrons participate in scattering processes, making it harder to enhance the single-particle charge gap with increasing $U$.
Figure 4: Mott gap at filling $\langle\hat{n}\rangle=1$ for SU($3$) $n$-OHK as a function of interaction strength $U$ for varying $n$. For the 8-OHK model, we use a $4\times 2$ unit cell.
@@ IV SU(3) antiferromagnetism
Apart from charge properties, we are also interested in generalized magnetic correlations in the SU($N$) models. As studied previously Feng et al. (2023); Ibarra-García-Padilla et al. (2023), we consider the equal-time structure factor for the same-spin or same-flavor species defined as
@@
where the transfer momentum $\bf q$ lies in the original full BZ. Constructing an OHK model based on a specific cluster scheme already selects a corresponding set of scattering momenta ${\bf q}_{s}$ (also within the original BZ) in the interaction term Mai et al. (2024). We focus on ${\bf q}={\bf q}_{s}$ since only the scattering of these momenta are considered. Then $S({\bf q})$ adopts a simple form (see supplement for a detailed derivation)
$S({\bf q})$ $=\frac{1}{nNV}\sum_{\alpha\beta{\bf k}\sigma}\left(\langle n_{\alpha{\bf k}\sigma}n_{\beta{\bf k}\sigma}\rangle-\langle n_{\alpha{\bf k}\sigma}\rangle\langle n_{\beta{\bf k}\sigma}\rangle\right)e^{i{\bf q}\cdot({\bf r}_{\alpha}-{\bf r}_{\beta})},$ (5)
@@
Here we calculate zero-temperature $S(\mathbf{q})$ of the SU($3$) 9 ($3\times 3$)-OHK model at filling $\langle\hat{n}\rangle=1$. The choice of $3\times 3$ clustering is to include the scattering of momentum $(2\pi/3,2\pi/3)$, which has been shown to be the leading magnetic correlations in SU(3) Hubbard model at a large $U$ Feng et al. (2023); Ibarra-García-Padilla et al. (2023). First, we determine the critical interaction strength for the metal-insulator transition in the 9-OHK model. In Fig 4, we fix $N=3$ and plot the $\langle\hat{n}\rangle=1$ Mott gap as a function of $U$ for varying orbital number $n=4,8,9$. We observed that the predicted $U_{c}$ is reasonably stable for the larger orbital numbers. For the 9-OHK, we find $U_{c}\simeq 5.4t$ which is in line with Hubbard model results obtained using DQMC ($U_{c}\simeq 6t$Ibarra-García-Padilla et al. (2023) and constrained-path anxiliary-field QMC ($U_{c}\simeq 5.5t$Feng et al. (2023).
@@
The $S(\mathbf{q})$ is shown in Fig. 5(a) as a function of $U>U_{c}$ at special momenta $\mathbf{q}=(0,0)$, $(2\pi/3,0)$ and $(2\pi/3,2\pi/3)$. We observe that the structure factor at $\mathbf{q}=(0,0)$ vanishes. For ${\bf q}=(2\pi/3,0)$, the structure factor is suppressed as the interaction strength increases. In contrast, the structure factor at ${\bf q}=(2\pi/3,2\pi/3)$ is enhanced. This result indicates that the OHK interaction favors $(2\pi/3,2\pi/3)$ antiferromagnetic (AFM) correlation, consistent with previous studies Feng et al. (2023).
@@
Another advantage of the OHK treatment over DQMC is the ease of obtaining dynamical correlations. Here we consider the dynamical same-spin structure factor
@@
$S({\bf q},\omega)=-\text{Im}\big{[}$ $\frac{1}{nNV}\sum_{\alpha\beta{\bf k}\sigma}\Big{(}\langle n_{\alpha{\bf k}\sigma}\frac{1}{\omega+i\eta+E_{0}-H}n_{\beta{\bf k}\sigma}\rangle$
@@
$+\langle n_{\alpha{\bf k}\sigma}\frac{1}{\omega-i\eta-E_{0}+H}n_{\beta{\bf k}\sigma}\rangle\Big{)}e^{i{\bf q}\cdot({\bf r}_{\alpha}-{\bf r}_{\beta})}\big{]},$ (6)
@@
at zero temperature. This quantity is computed accurately using the Lanczos method with the ground state obtained from exact diagonalization. The dynamical structure factor as a function $\omega$ at momenta ${\bf q}=(2\pi/3,0)$ and $(2\pi/3,2\pi/3)$ is shown in Fig. 5(b) and (c), respectively, for varying interaction strengths $U=6,8,10$. With increasing $U$, we observe the clear development of a low-energy pole ($\omega_{\rm peak}(2\pi/3,2\pi/3)\approx 0.4t$ for $U=10t$) for $S({\bf q}=(2\pi/3,2\pi/3),\omega)$. Contrastly, at ${\bf q}=(2\pi/3,0)$, the correlator shows a relatively weaker peak centred at a slightly higher frequency ($\omega_{\rm peak}(0,2\pi/3)\approx 0.7t$ for $U=10t$). This suggests the system has a propensity towards AFM magnetic ordering at wavevector $(2\pi/3,2\pi/3)$ with sub-leading fluctuations at ${\bf q}=(2\pi/3,0)$. This result agrees with the equal-time structure factor calculation.
@@
The development of $(2\pi/3,2\pi/3)$ AFM correlations at large $U$ under filling $\langle\hat{n}\rangle=1$ is consistent with results from SU($3$) Hubbard Feng et al. (2023); Ibarra-García-Padilla et al. (2023) and Heisenberg models Tóth et al. (2010); Bauer et al. (2012); Romen and Läuchli (2020). The OHK model exhibiting the same magnetic behaviour in the Mott insulating phase as the Hubbard model is another benchmark to support the validity of our approach.

### v3

@@ III.1 Mott insulator at ⟨$\langle\hat{n}\rangle=1$
⟨$\langle\hat{n}\rangle=\frac{1}{nV}\sum_{\begin{subarray}{c}\mathbf{k}\in{\rm rBZ}_{4},\\ \alpha,\sigma\end{subarray}}\langle n_{\alpha\mathbf{k}\sigma}\rangle,$ (8)
and compressibility $\chi\equiv\partial\langle\hat{n}\rangle/\partial\mu$ as a function of $\mu$ for the 4-OHK model in panels (a) and (c), and the corresponding DQMC results are shown in panels (b) and (d). Here, $V=L\times L$ is the number of momenta in rBZ4. We already find qualitative agreement between 4-OHK and the state-of-the-art DQMC simulations on the Hubbard model. This is expected given the rapid convergence of OHK to the Hubbard model. In particular, we observe the softening of the Mott gap as $N$ increases in both the density and compressibility. However, there are quantitative differences, for example, 4-OHK overestimates the magnitude of the Mott gap.
In Fig. 2(e,f), we present the benchmark for double occupancy $\mathcal{D}$, and its derivative ∂$\partial\mathcal{D}/\partial\langle\hat{n}\rangle$ as a function of density ⟨$\langle\hat{n}\rangle$. Double occupancy in the OHK model is defined as
$\mathcal{D}=\frac{1}{2nV}\sum_{\begin{subarray}{c}\mathbf{k}\in{\rm rBZ}_{4},\\ \alpha,\sigma\neq\sigma^{\prime}\end{subarray}}\langle n_{\alpha\mathbf{k}\sigma}n_{\alpha\mathbf{k}\sigma^{\prime}}\rangle.$ (9)
@@
This definition differs from that used in the SU($N$) Hubbard model. However, in both cases, $\mathcal{D}$ is proportional to the average interaction energy and the first derivative of the free energy with respect to $U$, thereby preserving this physical correspondence. In Fig. 2(e,f), the solid lines are the OHK results, while the markers denote DQMC-Hubbard results. The double occupancy from 4-OHK and DQMC-Hubbard agrees quantitatively. Moreover, 4-OHK shows consistent results at higher density ⟨$\langle\hat{n}\rangle>1$, whereas DQMC data becomes oscillatory due to the severe fermionic sign problem. As previously reported for SU($N$) Hubbard [37], the sharp jump in ∂$\partial\mathcal{D}/\partial\langle\hat{n}\rangle$ serves as a strong signature of the Mott insulating state persisting even in the high-temperature regime.
@@
From this benchmarking, we find that the 4-OHK model successfully captures the high-temperature $N$-dependent physics in the SU($N$) Hubbard model. Aside from the overestimate of the Mott gap, we find quantitative agreement between 4-OHK and DQMC-Hubbard results for the density ⟨$\langle\hat{n}\rangle$-dependence of physical quantities, such as $\mathcal{D}$. Additionally, the simplicity of 4-OHK makes it advantageous for exploring higher densities and lower temperatures (next section), where DQMC simulations are limited by the fermionic sign problem. These features establish OHK as a powerful tool for studying SU($N$) Hubbard physics beyond the reach of conventional numerical methods.
@@
Low temperature. In Fig. 3(a), we calculate the ⟨$\langle\hat{n}\rangle$ vs $\mu$ relation around ⟨$\langle\hat{n}\rangle=1$ at $U=8t$ and inverse temperature $\beta=200t^{-1}$ for the 4-OHK model with varying $N=2,3,4$. As expected from the high-temperature results in Fig. 2(a,b), we observe a sharp characteristic plateau, signaling the presence of a Mott charge gap. However, a surprising feature emerges: the Mott gap does not shrink monotonically with increasing $N$ despite appearing to do so at high temperatures in both 4-OHK (Fig. 2(a)) or DQMC simulations of the Hubbard model (Fig. 2(b,f)). Fig. 3(a) reveals the ordering $\Delta(N=3)>\Delta(N=4)>\Delta(N=2)$ for $U=8t$, deviating from the high-temperature behavior. This discrepancy arises because, at high temperatures, the number of excited states increases monotonically with $N$, leading to more pronounced thermal softening of the gap for larger $N$. By leveraging the 4-OHK model, we can explore both low- and high-temperature regimes, allowing us to disentangle intrinsic ground-state properties from thermal effects associated with excited states.
@@
Further, we calculate the charge gap at ⟨$\langle\hat{n}\rangle=1$ as a function of $U$ for various $N$ in the 4- and 8 (4$\times$2)-OHK models at zero temperature, as shown in Fig. 3(b) and (c), respectively. In the SU(2) case, a charge gap opens for any $U>0$ due to Fermi surface nesting at exactly ⟨$\langle\hat{n}_{\sigma}\rangle=1/2$ with ${\bf q}=(\pi,\pi)$ [27, 28, 33]. For $N>2$, such a nesting is absent at ⟨$\langle\hat{n}\rangle=1$, where ⟨$\langle\hat{n}_{\sigma}\rangle=1/N<1/2$. Consequently, a finite critical $U_{c}$ is required to open the gap, as indicated in Fig. 3(b). Moreover, we find that $U_{c}(N=3)>U_{c}(N=4)$, likely due to the particular Fermi surface geometry at $U=0$. This trend, $U_{c}(N=3)>U_{c}(N=4)$ is also observed in the 8-OHK (Fig. 3) with only quantitative modification. Once the gaps are open for both SU($3$) and SU($4$) cases, we observe $\Delta(N=3)>\Delta(N=4)$, consistent with Fig. 3(a). Extending our calculation to larger $N$, we find in Fig. 3(b) the results collapse for $N>n$. This can be understood from the structure of the OHK model, which consists of independent $n$-site Hubbard clusters at each $\mathbf{k}$ in the rBZn. When a charge gap opens at ⟨$\langle\hat{n}\rangle=1$, these $n$-site clusters are filled with exactly $n$ particles each. When $N>n$, these particles occupy different SU($N$) flavors to minimize the kinetic energy. Further increasing $N$ does change this configuration, meaning that $n$-OHK would not be able to capture any additional $N$-dependent physics beyond $N=n+1$. For this reason, for the band HK and 2-OHK, the ⟨$\langle\hat{n}\rangle=1$ results overlap for all SU($N>1$) and SU($N>2$), respectively (see the appendix). It is important to note that this phenomenon is specific to the filling ⟨$\langle\hat{n}\rangle=1$. As we will discuss in the next section, near half-filling ⟨$\langle\hat{n}\rangle=N/2$ when the charge gap opens, each Hubbard cluster in OHK contains $Nn/2$ particles. In this case, $n$-OHK retains the non-trivial SU($N$) physics for any $N$ as long as $n\geq 2$.
@@
Figure 4: Charge gap at half-filling (⟨$\langle\hat{n}\rangle=N/2$) as a function of interaction strength $U$ for SU($N$) 4-OHK model.
III.2 Mott insulator at ⟨$\langle\hat{n}\rangle=N/2$
As mentioned above, another generalization of the half-filled SU($2$) Mott insulator is the SU($N$) Mott insulator at half-filling ⟨$\langle\hat{n}\rangle=N/2$ for even $N$ (no half-filled charge gap exists for odd $N$). At this filling, the non-interacting Fermi surface for each flavor remains the same as in the SU($2$) case, thereby preserving the effect of Fermi surface nesting. Hence, we expect the charge gap opens for any finite $U>0$. This expectation is confirmed in Fig. 4 where we present the half-filled charge gap as a function of $U$ for the SU(N) 4-OHK model. For a fixed interaction strength, the charge gap decreases monotonically with increasing $N$. As the number of flavors increases, more electrons participate in scattering processes, making it harder to enhance the single-particle charge gap with increasing $U$.
Figure 5: Mott gap at filling ⟨$\langle\hat{n}\rangle=1$ for SU($3$) $n$-OHK as a function of interaction strength $U$ for varying $n$. For the 8-OHK model, we use a $4\times 2$ unit cell.
@@ IV SU(3) antiferromagnetism
Apart from charge properties, we are also interested in generalized magnetic correlations in the SU($N$) models. As previously studied [27, 28], we consider the equal-time structure factor for the same-spin or same-flavor species defined as
@@
where the transfer momentum $\bf q$ lies in the original full BZ. Constructing an OHK model based on a specific cluster scheme already selects a corresponding set of scattering momenta ${\bf q}_{s}$ (also within the original BZ) in the interaction term [33]. We focus on ${\bf q}={\bf q}_{s}$ since only the scattering of these momenta are considered. Then $S({\bf q})$ adopts a simple form (see the appendix for a detailed derivation)
$S({\bf q})=\frac{1}{nNV}\sum_{\begin{subarray}{c}\mathbf{k}\in{\rm rBZ}_{n},\\ \alpha\alpha^{\prime},\sigma\end{subarray}}$ (⟨$\big{(}\langle n_{\alpha{\bf k}\sigma}n_{\alpha^{\prime}{\bf k}\sigma}\rangle$
−⟨$-\langle n_{\alpha{\bf k}\sigma}\rangle\langle n_{\alpha^{\prime}{\bf k}\sigma}\rangle\big{)}e^{i{\bf q}\cdot({\bf r}_{\alpha}-{\bf r}_{\alpha^{\prime}})},$ (11)
@@
Here we calculate zero-temperature $S(\mathbf{q})$ of the SU($3$) 9-OHK model ($3\times 3$ unit cell) at filling ⟨$\langle\hat{n}\rangle=1$. The choice of $3\times 3$ clustering is to include the scattering of momentum $(2\pi/3,2\pi/3)$, which has been shown to be the leading magnetic correlations in the SU(3) Hubbard model at a large $U$ [27, 28]. First, we determine the critical interaction strength for the metal-insulator transition in the 9-OHK model. In Fig 5, we fix $N=3$ and plot the ⟨$\langle\hat{n}\rangle=1$ Mott gap as a function of $U$ for varying orbital number $n=4,8,9$. We observed that the predicted $U_{c}$ is reasonably stable for the larger orbital numbers. For the 9-OHK, we find $U_{c}\simeq 5.4t$, which is in line with the Hubbard model results obtained using DQMC ($U_{c}\simeq 6t$) [28] and constrained-path auxiliary-field QMC ($U_{c}\simeq 5.5t$) [27].
@@
The $S(\mathbf{q})$ is shown in Fig. 6(a) as a function of $U>U_{c}$ at special momenta $\mathbf{q}=(0,0)$, $(2\pi/3,0)$ and $(2\pi/3,2\pi/3)$. We observe that the structure factor at $\mathbf{q}=(0,0)$ vanishes. For ${\bf q}=(2\pi/3,0)$, the structure factor is suppressed as the interaction strength increases. In contrast, the structure factor at ${\bf q}=(2\pi/3,2\pi/3)$ is enhanced. This result indicates that the OHK interaction favors $(2\pi/3,2\pi/3)$ antiferromagnetic (AFM) correlation, consistent with previous studies [27].
@@
Another advantage of the OHK treatment over DQMC is the ease of obtaining dynamical correlations (the local spectral function of OHK is shown in the appendix and compared to dynamical mean-field theory [39]). Here we consider the dynamical same-spin structure factor
@@
$S({\bf q},\omega)=-$ $\text{Im}\Bigg{[}\frac{1}{nNV}\sum_{\begin{subarray}{c}\mathbf{k}\in{\rm rBZ}_{n},\\ \alpha\alpha^{\prime},\sigma\end{subarray}}\Big{(}\langle n_{\alpha{\bf k}\sigma}\frac{1}{\omega+i\eta+E_{0}-H}n_{\alpha^{\prime}{\bf k}\sigma}\rangle$
@@
+⟨$+\langle n_{\alpha^{\prime}{\bf k}\sigma}\frac{1}{\omega-i\eta-E_{0}+H}n_{\alpha^{\prime}{\bf k}\sigma}\rangle\Big{)}e^{i{\bf q}\cdot({\bf r}_{\alpha}-{\bf r}_{\alpha^{\prime}})}\Bigg{]},$ (12)
@@
at zero temperature. This quantity is computed accurately using the Lanczos method with the ground state obtained from exact diagonalization. The dynamical structure factor as a function $\omega$ at momenta ${\bf q}=(2\pi/3,0)$ and $(2\pi/3,2\pi/3)$ is shown in Fig. 6(b) and (c), respectively, for varying interaction strengths $U=6,8,10$. With increasing $U$, we observe the clear development of a low-energy pole ($\omega_{\rm peak}(2\pi/3,2\pi/3)\approx 0.4t$ for $U=10t$) for $S({\bf q}=(2\pi/3,2\pi/3),\omega)$. In contrast, at ${\bf q}=(2\pi/3,0)$, the correlator shows a relatively weaker peak centered at a slightly higher frequency ($\omega_{\rm peak}(0,2\pi/3)\approx 0.7t$ for $U=10t$). This suggests the system has a propensity towards AFM magnetic ordering at wavevector $(2\pi/3,2\pi/3)$ with sub-leading fluctuations at ${\bf q}=(2\pi/3,0)$. This result agrees with the equal-time structure factor calculation.
@@
The development of $(2\pi/3,2\pi/3)$ AFM correlations at large $U$ under filling ⟨$\langle\hat{n}\rangle=1$ is consistent with results from SU($3$) Hubbard [27, 28] and Heisenberg models [18, 19, 21]. The OHK model exhibiting the same magnetic behavior in the Mott insulating phase as the Hubbard model is another benchmark to support the validity of our approach.